### v1

@@ Abstract
State space models (SSMs) with efficient hardware-aware architectures, namely Mamba, have recently demonstrated significant promise in long-sequence modeling. Researchers are now exploring how to adapt Mamba for computer vision tasks. This paper is the first comprehensive survey aiming to provide an in-depth analysis of Mamba models in the field of computer vision. It begins by exploring the foundational concepts contributing to Mamba’s success, including the state space model framework, selection mechanisms, and hardware-aware design. Next, we review these vision mamba models by categorizing them into foundational mamba and using other techniques to make them more refined. We further delve into the widespread applications of transformers in vision tasks, which include their use as a backbone and in various levels of vision processing. This encompasses High/Mid-level vision (e.g., Object detection, Segmentation, Video classification, etc.), Low-level vision (e.g., Image super-resolution, Image restoration, Visual generation, etc.), Medical image (e.g., 2D / 3D Segmentation, Classification, and Deformable image registration, etc.) and Remote Sensing Image. We hope this endeavor will spark additional interest within the community to address current challenges and further apply Mamba models in computer vision.
@@
Mamba; Computer vision; State space model.
[…]
@@ 1 Introduction
Deep Neural Networks (DNNs) have demonstrated remarkable performance across various artificial intelligence (AI) tasks, with the fundamental architecture playing a crucial role in determining the model’s capabilities. Traditional neural networks typically comprise Multi-Layer Perceptron (MLP) or Fully Connected (FC) layers rosenblatt1957perceptron; rosenblatt1962principles. Convolutional neural networks (CNNs) lecun1998gradient; krizhevsky2012imagenet introduce convolutional and pooling layers, which are particularly effective for processing shift-invariant data like images. Recurrent neural networks (RNNs) hochreiter1997long; bahdanau2014neural utilize recurrent cells to handle sequential or time series data. To address the issue of CNN, RNN, and GNN models only capturing local relationships, the Transformer parikh2016decomposable; vaswani2017attention; devlin2018bert, introduced in 2017, excels at learning long-distance feature representations. Transformers primarily rely on attention-based attention mechanisms, e.g. self-attention and cross-attention, to extract intrinsic features and improve their representation capability. Pre-trained massive transformer-based models, such as GPT-3 brown2020language, deliver robust performance across various NLP datasets, excelling in natural language understanding and generation tasks. The remarkable performance of transformer-based models has propelled their widespread adoption in vision applications. The core of transformer models is their exceptional skill in capturing long-range dependencies and maximizing the use of large datasets. The feature extraction module is the main component of vision transformer architectures. It processes data using a sequence of self-attention blocks, significantly improving its capacity to analyze images.
@@
As SSMs have evolved, a new class of selective state space models, termed Mamba gu2023mamba. It has advanced the modeling of discrete data, such as text, with state-space models (SSMs) through two key improvements. Firstly, it features an input-dependent mechanism that adjusts SSM parameters dynamically, enhancing information filtering. Secondly, Mamba uses a hardware-aware algorithm that processes data linearly with sequence length, boosting computational speed on modern systems. Inspired by Mamba’s achievements in language modeling, several initiatives are now aiming to adapt this success to the field of vision. Several studies have explored its integration with Mixture-of-Experts (MoE) techniques, as evidenced by works like Jamba lieber2024jamba, MoE-Mamba pioro2024moe, and BlackMamba anthony2024blackmamba, outperformed the state-of-the-art architecture Transformer-MoE with fewer training steps. As illustrated in Fig. 1(b), since the release of Mamba in December 2023, the number of research papers focusing on Mamba in the vision domain has rapidly increased, reaching a peak in March 2024. This trend suggests that Mamba is emerging as a prominent research area in vision, potentially offering a viable alternative to Transformers. Therefore, A review of current related works is necessary and timely to provide a detailed overview of this new methodology in this evolving field.
@@ 2 Formulation of Mamba
Mamba gu2023mamba is initially introduced in the domain of natural language processing. As shown in Fig. 2, the original Mamba Block integrates a Gated MLP into the State Space Model (SSM) architecture of H3 fu2022hungry, utilizing an SSM sandwiched between two gated connections alongside a standard local convolution. For $\sigma$ SiLU hendrycks2016gaussian or Swish activation function Ramachandran2017SwishAS is used. The Mamba architecture consists of repeating of Mamba block interleaved with standard normalization and residual connections. An optional normalization layer (LayerNorm chosen by original Mamba) is used in a similar location as RetNet sun2023retentive.
Figure 2: Mamba Block gu2023mamba.
@@ 2.1 State Space Models (SSMs)
Consider a structured state space model (SSM) that maps one-dimensional sequence $x(t)\in^{L}$ to $y(t)\in^{L}$ through a hidden state $h(t)\in^{N}$. With the evolution parameter $\mathbf{A}\in^{N\times N}$ and the projection parameters $\mathbf{B}\in^{N\times 1}$, $\mathbf{C}\in^{1\times N}$, such a model is formulated as linear ordinary differential equations
@@
where $L$ is the length of the input sequence $\mathbf{x}$, $\overline{\mathbf{K}}\in^{L}$ serves as the kernel of the SSMs and * represents the convolution operation.
@@
SSMs often serve as independent sequence transformations that can be integrated into end-to-end neural network architectures. Here we introduce several fundamental architectures. Linear attention katharopoulos2020transformers approximates self-attention with a recurrence mechanism as a simplified form of linear SSM. H3 fu2022hungry, as illustrated in Fig. 2, places an SSM between two gated connections and inserts a standard local convolution before it. Following H3, Hyena poli2023hyena, replaces the SSM layer with an MLP-parameterized global convolution romero2021ckconv. RetNet sun2307retentive introduces an extra gate and uses simpler SSM. RetNet enables an alternative parallelizable computation path and employs a variant of multi-head attention (MHA) instead of convolutions. Inspired by attention-free Transformer zhai2021attention, the recent RNN design RWKV peng2023rwkv, can be interpreted as the ratio of two SSMs due to its primary "WKV" mechanism involving Linear Time Invariance (LTI) recurrences.
@@
before (2,3), so that parameters $\mathbf{B}\in^{B\times L\times N}$, $\mathbf{C}^{B\times L\times N}$ and $\Delta^{B\times L\times D}$ are dependent on the input sequence $\mathbf{x}\in^{B\times L\times D}$, where $B$ represents the batch size, and $D$ represents number of channels. Normally, $S_{B}$ and $S_{C}$ are linear parameterized projections to dimension $N$, i.e. $Linear_{N}(\cdot)$, while $S_{\Delta}(\mathbf{x})=Broadcast_{D}(Linear_{1}(\mathbf{x}))$ and $\tau_{\Delta}=softplus$. The choice of $S_{\Delta}$ and $\tau_{\Delta}$ is due to a connection to RNN gating mechanisms explained later.
@@ 2.2 Other Key Concepts in Mamba
The connection between RNN gating and the discretization of continuous-time systems is well-established tallec2018can. The classical gating mechanism of RNNs is an instance of the selection mechanism for SSMs. When $N=1,\mathbf{A}=-1,\mathbf{B}=1,S_{\Delta}=Linear(\mathbf{x})$ and $\tau_{\Delta}=softplus$, then the selective SSM recurrence takes the form
@@ 3 Mamba for Vision
In this section, we present efforts aimed at enabling Mamba to tackle vision-related tasks while enhancing its efficiency and performance. Initially, we delve into two foundational works: Vision Mamba zhu2024vision and VMamaba liu2024vmamba. These works introduced the ViM block and VSS block, respectively, serving as the foundation for subsequent research endeavors. Subsequently, we explore additional works focused on refining the Mamba architecture as a backbone for vision-related tasks. Lastly, we discuss the work of integrating Mamba with other architectures such as convolution, recurrence, and attention.
@@
ViM block zhu2024vision, sometimes also mentioned as Bidirectional Mamba block, annotates image sequences with position embeddings and condenses visual representations using bidirectional state space models. It processes input both forward and backward, employing one-dimensional convolution for each direction as shown in (a) Fig. 4. The Softplus function ensures non-negative $\Delta$. Forward and backward $\mathbf{y}$ are computed via the state space model described in equations (2) and (3), and then combined through SiLU gating to produce the output token sequence as (a) Fig. 3.
@@
The Visual State Space (VSS) block  liu2024vmamba incorporates the pivotal state space model operation. It begins by directing the input through a depth-wise convolution layer, followed by a SiLU activation function, and then through the state space model outlined in equations (2) and (3) employing an approximate $\overline{\mathbf{B}}$. Afterwards, the output of the state space model is subjected to layer normalization before being amalgamated with the output of other information streams as (b) Fig. 3. To address the encountered direction-sensitive issue, they introduce the Cross-Scan Module (CSM) to traverse the spatial domain and convert any non-causal visual image into order patch sequences, as shown in (b) Fig. 4. They refine the approximation of $\overline{\mathbf{B}}$ using the first-order Taylor series $\overline{\mathbf{B}}=(\Delta\mathbf{A})^{-1}(\exp(\Delta\mathbf{A})-\mathbf{I})\cdot\Delta\mathbf{B}\thickapprox(\Delta\mathbf{A})(\Delta\mathbf{A})^{-1}\Delta\mathbf{B}=\Delta\mathbf{B}$.
@@ 3.2 Pure Mamba
Inspired by the vision transformer architecture, Vision Mamba zhu2024vision replaces the transformer encoder with a vision mamba encoder based on ViM blocks while retaining the remainder of the process. This involves converting the two-dimensional image into flattened patches, followed by linear projection of these patches into vectors and the addition of position embeddings. A class token represents the entire patch sequence, and subsequent steps involve normalization layers and a MLP layer to derive the final predictions.
LocalMamba huang2024localmamba is built based on Vim block and it introduces a novel scanning methodology that includes localized scanning within distinct windows to capture detailed local information in conjunction with global context. Plus, LocalMamba searches scanning directions across different network layers to identify and apply the most effective scanning combinations. They propose two variants, i.e. with plain and hierarchical structures. They proposed their LocalVim Block including four scanning directions (cf. (d) Fig. 4): vim scanning and partitions tokens into distinct windows along with their flipped counterparts facilitating scanning from tail to head, state space module, and spatial and channel attention module (SCAttn).
@@
VMamba liu2024vmamba has four stages after partitioning the input image into patches as Vision Mamba. VMamba stacks several VSS blocks on the feature map with resolution $\frac{H}{4}\times\frac{W}{4}$ as Stage 1. In Stage 2, before more VSS blocks involving, the feature map in Stage 1 goes through a patch merge operation for down-sampling in order to build hierarchical representations, resulting in an output resolution of $\frac{H}{8}\times\frac{W}{8}$. While Stage 3 and Stage 4 are the repetition of Stage 1 and Stage 2 with resolutions of $\frac{H}{16}\times\frac{W}{16}$ and $\frac{H}{32}\times\frac{W}{32}$.
Based on VSS block, PlainMamba block yang2024plainmamba enhances its capability to learn features from two-dimensional images through two main mechanisms: (i) employing a continuous 2D scanning process to improve spatial continuity, ensuring tokens in the scanning sequence are adjacent, as illustrated in (c) Fig. 4, and (ii) incorporating direction-aware updating to enable the model to discern spatial relations among tokens by encoding directional information. PlainMamba improves the spatial discontinuity when moving to a new row/column in the 2D scanning mechanisms of Vim and VMamba by continuing the scanning with a reversed direction until it reaches the final vision token of the image. Furthermore, PlainMamba eliminates the need for special tokens.
Within lightweight model designs, EfficientVMamba pei2024efficientvmamba improves the capabilities of VMamba with an atrous-based selective scan approach, i.e. Efficient 2D Scanning (ES2D). Instead of scanning all patches from various directions and increasing the total number of patches, ES2D adopts a strategy of scanning forward vertically and horizontally while skipping patches and maintaining the number of patches unchanged, as shown in (e) Fig. 4. Their Efficient Visual State Space (EVSS) block comprises a convolutional branch for local features, uses ES2D as the SSM branch for global features, and all branches end through a squeeze-excitation block. They employ EVSS blocks for both Stage 1 and Stage 2, while opting for Inverted Residual blocks in Stage 3 and Stage 4 to enhance the capture of global representations.
@@
As a part of multi-dimensional data, existing models for multi-dimensional data also work for visual-related tasks but often lack the capacity to facilitate inter- and intra-dimension communication or data-independent. The MambaMixer block behrouz2024mambamixer introduces a dual selection mechanism spanning across tokens and channels. It then links the sequential selective mixers through a weighted averaging mechanism, enabling layers to directly access input and output from various layers. Mamba-ND li2024mamba expands the application of the SSM to higher dimensions by alternating sequence wandering across layers. Utilizing a similar scanning strategy as VMamba in the 2D scenario, it extends this approach to 3D. Additionally, they advocate for the use of multi-head SSM as an analogue to multi-head attention. In response to the inefficiencies and performance challenges encountered by traditional transformers in image and time series processing, a new architecture named Simplified Mamba-based Architecture, SiMBA patro2024simba is proposed to incorporate the Mamba block for sequence modeling and Einstein FFT (EinFFT) for channel modeling, with the goal of enhancing the stability and efficiency of the model in handling image and time series tasks. The Mamba block proves effective at processing long sequence data, while EinFFT represents a novel channel modeling technique. Experimental results demonstrate that SiMBA surpasses existing State Space Models and transformers across multiple benchmark tests.
@@
(a) BiDirectional (b) Cross-Scan liu2024vmamba (c) Continuous 2D (d) Local Scan huang2024localmamba
Scan zhu2024vision    Scanning yang2024plainmamba
@@
(e) Efficient 2D Scanning (ES2D) pei2024efficientvmamba (f) Zigzag Scan hu2024zigma
@@
(g) Omnidirectional Selective Scan shi2024vmambair (h) 3D BiDirectional Scan li2024videomamba
@@
(i) Hierarchical Scan zhang2024motion (j) Spatiotemporal Selective Scan yang2024vivim
(k) Multi-Path Scan chen2024rsmamba
@@
As an important element of Mamba, scanning mechanisms not only help in efficiency but also provide information in the scenario of visual-related tasks. We summarize the usage of different scanning mechanisms in existing works as Table. 1. Cross-Scan liu2024vmamba and BiDirectional Scan zhu2024vision stand out as the most widely adopted scanning mechanisms. However, various other scanning mechanisms serve specific purposes. For example, 3D BiDirectional Scan li2024videomamba and Spatiotemporal Selective Scan yang2024vivim are tailored for video inputs. Local Scan huang2024localmamba focuses on gathering local information, while ES2D pei2024efficientvmamba prioritizes efficiency.
@@
BiDirectional Scan zhu2024vision Vision Mamba zhu2024vision,Motion Mamba zhang2024motion
HARMamba li2024harmamba,MMA cheng2024activating,VL-Mambaqiao2024vl
Video Mamba Suite chen2024video,Point Mamba liu2024point
LMa-UNet wang2024large
Motion-Guided Dual-Camera Tracker zhang2024motion2
Cross-Scan liu2024vmamba VMamba liu2024vmamba,VL-Mambaqiao2024vl,VMRNN tang2024vmrnn
RES-VMAMBA chen2024res,Sigma wan2024sigma,ReMamber yang2024remamber
Mamba-UNet wang2024mamba,Semi-Mamba-UNet ma2024semi
VMambaMorph wang2024vmambamorph,ChangeMamba chen2024changemamba
H-vmunet wu2024h,MambaMIR huang2024mambamir,MambaIR guo2024mambair
Serpent shahab2024serpent,Mamba-HUNet shahriar2024integrating,TM-UNet tang2024rotate
Swin-UMamba liu2024swin,UltraLight VM-UNet wu2024ultralight
VM-UNet ruan2024vm,VM-UNET-V2  zhang2024vm
MedMamba yue2024medmamba,MIM-ISTD chen2024mim,RS3Mamba ma2024rs3mamba
Continuous 2D Scanning yang2024plainmamba PlainMamba yang2024plainmamba
Local Scan huang2024localmamba LocalMamba huang2024localmamba,FreqMamba zhen2024freqmamba
Efficient 2D Scanning (ES2D) pei2024efficientvmamba EfficientVMamba pei2024efficientvmamba
Zigzag Scan hu2024zigma ZigMa hu2024zigma
Omnidirectional Selective Scan shi2024vmambair VmambaIR shi2024vmambair,RS-Mamba zhao2024rs
3D BiDirectional Scan li2024videomamba VideoMamba li2024videomamba
Hierarchical Scan zhang2024motion Motion Mamba zhang2024motion
Spatiotemporal Selective Scan yang2024vivim Vivim yang2024vivim
Multi-Path Scan chen2024rsmamba RSMamba chen2024rsmamba

### v2

@@ Abstract
State space models (SSMs) with selection mechanisms and hardware-aware architectures, namely Mamba, have recently demonstrated significant promise in long-sequence modeling. Since the self-attention mechanism in transformers has quadratic complexity with image size and increasing computational demands, the researchers are now exploring how to adapt Mamba for computer vision tasks. This paper is the first comprehensive survey aiming to provide an in-depth analysis of Mamba models in the field of computer vision. It begins by exploring the foundational concepts contributing to Mamba’s success, including the state space model framework, selection mechanisms, and hardware-aware design. Next, we review these vision mamba models by categorizing them into foundational ones and enhancing them with techniques such as convolution, recurrence, and attention to improve their sophistication. We further delve into the widespread applications of Mamba in vision tasks, which include their use as a backbone in various levels of vision processing. This encompasses general visual tasks, Medical visual tasks (e.g., 2D / 3D segmentation, classification, and image registration, etc.), and Remote Sensing visual tasks. We specially introduce general visual tasks from two levels: High/Mid-level vision (e.g., Object detection, Segmentation, Video classification, etc.) and Low-level vision (e.g., Image super-resolution, Image restoration, Visual generation, etc.). We hope this endeavor will spark additional interest within the community to address current challenges and further apply Mamba models in computer vision.
@@
Mamba Computer vision State space model Application.
@@ 1 Introduction
Deep Neural Networks (DNNs) have demonstrated remarkable performance across various artificial intelligence (AI) tasks, with the fundamental architecture playing a crucial role in determining the model’s capabilities. Traditional neural networks typically comprise Multi-Layer Perceptron (MLP) or Fully Connected (FC) layers [1, 2]. Convolutional neural networks (CNNs) [3, 4] introduce convolutional and pooling layers, which are particularly effective for processing shift-invariant data like images. Recurrent neural networks (RNNs) [5, 6] utilize recurrent cells to handle sequential or time series data. To address the issue of CNN, RNN, and GNN models only capturing local relationships, the Transformer [7, 8, 9], introduced in 2017, excels at learning long-distance feature representations. Transformers primarily rely on attention-based attention mechanisms, e.g. self-attention and cross-attention, to extract intrinsic features and improve their representation capability. Pre-trained massive transformer-based models, such as GPT-3 [10], deliver robust performance across various NLP datasets, excelling in natural language understanding and generation tasks. The remarkable performance of transformer-based models has propelled their widespread adoption in vision applications. The core of transformer models is their exceptional skill in capturing long-range dependencies and maximizing the use of large datasets. The feature extraction module is the main component of vision transformer architectures. It processes data using a sequence of self-attention blocks, significantly improving its capacity to analyze images.
@@
As SSMs have evolved, a new class of selective state space models, termed Mamba [11]. It has advanced the modeling of discrete data, such as text, with state-space models (SSMs) through two key improvements. Firstly, it features an input-dependent mechanism that adjusts SSM parameters dynamically, enhancing information filtering. Secondly, Mamba uses a hardware-aware algorithm that processes data linearly with sequence length, boosting computational speed on modern systems. Inspired by Mamba’s achievements in language modeling, several initiatives are now aiming to adapt this success to the field of vision. Several studies have explored its integration with Mixture-of-Experts (MoE) techniques, as evidenced by works like Jamba [12], MoE-Mamba [13], and BlackMamba [14], outperformed the state-of-the-art architecture Transformer-MoE with fewer training steps. As illustrated in Fig. 1(b), since the release of Mamba in December 2023, the number of research papers focusing on Mamba in the vision domain has rapidly increased, reaching a peak in March 2024. This trend suggests that Mamba is emerging as a prominent research area in vision, potentially offering a viable alternative to Transformers. Therefore, A review of current related works is necessary and timely to provide a detailed overview of this new methodology in this evolving field.
@@ 2 Formulation of Mamba
Mamba [11] is initially introduced in the domain of natural language processing. As shown in Fig. 2, the original Mamba Block integrates a Gated MLP into the State Space Model (SSM) architecture of H3 [15], utilizing an SSM sandwiched between two gated connections alongside a standard local convolution. For $\sigma$ SiLU [16] or Swish activation function [17] is used. The Mamba architecture consists of repeating of Mamba block interleaved with standard normalization and residual connections. An optional normalization layer (LayerNorm chosen by original Mamba) is used in a similar location as RetNet [18].
Figure 2: Mamba Block [11].
@@ 2.1 State Space Models (SSMs)
Consider a structured state space model (SSM) that maps one-dimensional sequence $x(t)\in\mathbb{R}^{L}$ to $y(t)\in\mathbb{R}^{L}$ through a hidden state $h(t)\in\mathbb{R}^{N}$. With the evolution parameter $\mathbf{A}\in\mathbb{R}^{N\times N}$ and the projection parameters $\mathbf{B}\in\mathbb{R}^{N\times 1}$, $\mathbf{C}\in\mathbb{R}^{1\times N}$, such a model is formulated as linear ordinary differential equations
@@
where $L$ is the length of the input sequence $\mathbf{x}$, $\overline{\mathbf{K}}\in\mathbb{R}^{L}$ serves as the kernel of the SSMs and * represents the convolution operation.
@@
SSMs often serve as independent sequence transformations that can be integrated into end-to-end neural network architectures. Here we introduce several fundamental architectures. Linear attention [19] approximates self-attention with a recurrence mechanism as a simplified form of linear SSM. H3 [15], as illustrated in Fig. 2, places an SSM between two gated connections and inserts a standard local convolution before it. Following H3, Hyena [20], replaces the SSM layer with an MLP-parameterized global convolution [21]. RetNet [22] introduces an extra gate and uses simpler SSM. RetNet enables an alternative parallelizable computation path and employs a variant of multi-head attention (MHA) instead of convolutions. Inspired by attention-free Transformer [23], the recent RNN design RWKV [24], can be interpreted as the ratio of two SSMs due to its primary "WKV" mechanism involving Linear Time Invariance (LTI) recurrences.
@@
before (2,3), so that parameters $\mathbf{B}\in\mathbb{R}^{B\times L\times N}$, $\mathbf{C}^{B\times L\times N}$ and $\Delta^{B\times L\times D}$ are dependent on the input sequence $\mathbf{x}\in\mathbb{R}^{B\times L\times D}$, where $B$ represents the batch size, and $D$ represents number of channels. Normally, $S_{B}$ and $S_{C}$ are linear parameterized projections to dimension $N$, i.e. $Linear_{N}(\cdot)$, while $S_{\Delta}(\mathbf{x})=Broadcast_{D}(Linear_{1}(\mathbf{x}))$ and $\tau_{\Delta}=softplus$. The choice of $S_{\Delta}$ and $\tau_{\Delta}$ is due to a connection to RNN gating mechanisms explained later.
@@ 2.2 Other Key Concepts in Mamba
The connection between RNN gating and the discretization of continuous-time systems is well-established [25]. The classical gating mechanism of RNNs is an instance of the selection mechanism for SSMs. When $N=1,\mathbf{A}=-1,\mathbf{B}=1,S_{\Delta}=Linear(\mathbf{x})$ and $\tau_{\Delta}=softplus$, then the selective SSM recurrence takes the form
@@ 3 Mamba for Vision
In this section, we present efforts aimed at enabling Mamba to tackle vision-related tasks while enhancing its efficiency and performance. Initially, we delve into two foundational works: Vision Mamba [26] and VMamaba [27]. These works introduced the ViM block and VSS block, respectively, serving as the foundation for subsequent research endeavors. Subsequently, we explore additional works focused on refining the Mamba architecture as a backbone for vision-related tasks. Lastly, we discuss the work of integrating Mamba with other architectures such as convolution, recurrence, and attention.
@@
ViM block [26], sometimes also mentioned as Bidirectional Mamba block, annotates image sequences with position embeddings and condenses visual representations using bidirectional state space models. It processes input both forward and backward, employing one-dimensional convolution for each direction as shown in (a) Fig. 4. The Softplus function ensures non-negative $\Delta$. Forward and backward $\mathbf{y}$ are computed via the state space model described in equations (2) and (3), and then combined through SiLU gating to produce the output token sequence as (a) Fig. 3.
@@
The Visual State Space (VSS) block  [27] incorporates the pivotal state space model operation. It begins by directing the input through a depth-wise convolution layer, followed by a SiLU activation function, and then through the state space model outlined in equations (2) and (3) employing an approximate $\overline{\mathbf{B}}$. Afterwards, the output of the state space model is subjected to layer normalization before being amalgamated with the output of other information streams as (b) Fig. 3. To address the encountered direction-sensitive issue, they introduce the Cross-Scan Module (CSM) to traverse the spatial domain and convert any non-causal visual image into order patch sequences, as shown in (b) Fig. 4. They refine the approximation of $\overline{\mathbf{B}}$ using the first-order Taylor series $\overline{\mathbf{B}}=(\Delta\mathbf{A})^{-1}(\exp(\Delta\mathbf{A})-\mathbf{I})\cdot\Delta\mathbf{B}\thickapprox(\Delta\mathbf{A})(\Delta\mathbf{A})^{-1}\Delta\mathbf{B}=\Delta\mathbf{B}$.
@@ 3.2 Pure Mamba
Inspired by the vision transformer architecture, Vision Mamba [26] replaces the transformer encoder with a vision mamba encoder based on ViM blocks while retaining the remainder of the process. This involves converting the two-dimensional image into flattened patches, followed by linear projection of these patches into vectors and the addition of position embeddings. A class token represents the entire patch sequence, and subsequent steps involve normalization layers and a MLP layer to derive the final predictions.
LocalMamba [28] is built based on Vim block and it introduces a novel scanning methodology that includes localized scanning within distinct windows to capture detailed local information in conjunction with global context. Plus, LocalMamba searches scanning directions across different network layers to identify and apply the most effective scanning combinations. They propose two variants, i.e. with plain and hierarchical structures. They proposed their LocalVim Block including four scanning directions (cf. (d) Fig. 4): vim scanning and partitions tokens into distinct windows along with their flipped counterparts facilitating scanning from tail to head, state space module, and spatial and channel attention module (SCAttn).
@@
VMamba [27] has four stages after partitioning the input image into patches as Vision Mamba. VMamba stacks several VSS blocks on the feature map with resolution $\frac{H}{4}\times\frac{W}{4}$ as Stage 1. In Stage 2, before more VSS blocks involving, the feature map in Stage 1 goes through a patch merge operation for down-sampling in order to build hierarchical representations, resulting in an output resolution of $\frac{H}{8}\times\frac{W}{8}$. While Stage 3 and Stage 4 are the repetition of Stage 1 and Stage 2 with resolutions of $\frac{H}{16}\times\frac{W}{16}$ and $\frac{H}{32}\times\frac{W}{32}$.
Based on VSS block, PlainMamba block [29] enhances its capability to learn features from two-dimensional images through two main mechanisms: (i) employing a continuous 2D scanning process to improve spatial continuity, ensuring tokens in the scanning sequence are adjacent, as illustrated in (c) Fig. 4, and (ii) incorporating direction-aware updating to enable the model to discern spatial relations among tokens by encoding directional information. PlainMamba improves the spatial discontinuity when moving to a new row/column in the 2D scanning mechanisms of Vim and VMamba by continuing the scanning with a reversed direction until it reaches the final vision token of the image. Furthermore, PlainMamba eliminates the need for special tokens.
Within lightweight model designs, EfficientVMamba [30] improves the capabilities of VMamba with an atrous-based selective scan approach, i.e. Efficient 2D Scanning (ES2D). Instead of scanning all patches from various directions and increasing the total number of patches, ES2D adopts a strategy of scanning forward vertically and horizontally while skipping patches and maintaining the number of patches unchanged, as shown in (e) Fig. 4. Their Efficient Visual State Space (EVSS) block comprises a convolutional branch for local features, uses ES2D as the SSM branch for global features, and all branches end through a squeeze-excitation block. They employ EVSS blocks for both Stage 1 and Stage 2, while opting for Inverted Residual blocks in Stage 3 and Stage 4 to enhance the capture of global representations.
@@
As a part of multi-dimensional data, existing models for multi-dimensional data also work for visual-related tasks but often lack the capacity to facilitate inter- and intra-dimension communication or data-independent. The MambaMixer block [31] introduces a dual selection mechanism spanning across tokens and channels. It then links the sequential selective mixers through a weighted averaging mechanism, enabling layers to directly access input and output from various layers. Mamba-ND [32] expands the application of the SSM to higher dimensions by alternating sequence wandering across layers. Utilizing a similar scanning strategy as VMamba in the 2D scenario, it extends this approach to 3D. Additionally, they advocate for the use of multi-head SSM as an analogue to multi-head attention. In response to the inefficiencies and performance challenges encountered by traditional transformers in image and time series processing, a new architecture named Simplified Mamba-based Architecture, SiMBA [33] is proposed to incorporate the Mamba block for sequence modeling and Einstein FFT (EinFFT) for channel modeling, with the goal of enhancing the stability and efficiency of the model in handling image and time series tasks. The Mamba block proves effective at processing long sequence data, while EinFFT represents a novel channel modeling technique. Experimental results demonstrate that SiMBA surpasses existing State Space Models and transformers across multiple benchmark tests.
@@
(a) BiDirectional (b) Cross-Scan [27] (c) Continuous 2D (d) Local Scan [28]
Scan [26]    Scanning [29]
@@
(e) Efficient 2D Scanning (ES2D) [30] (f) Zigzag Scan [34]
@@
(g) Omnidirectional Selective Scan [35] (h) 3D BiDirectional Scan [36]
@@
(i) Hierarchical Scan [37] (j) Spatiotemporal Selective Scan [38]
(k) Multi-Path Scan [39]
@@
As an important element of Mamba, scanning mechanisms not only help in efficiency but also provide information in the scenario of visual-related tasks. We summarize the usage of different scanning mechanisms in existing works as Table. 1. Cross-Scan [27] and BiDirectional Scan [26] stand out as the most widely adopted scanning mechanisms. However, various other scanning mechanisms serve specific purposes. For example, 3D BiDirectional Scan [36] and Spatiotemporal Selective Scan [38] are tailored for video inputs. Local Scan [28] focuses on gathering local information, while ES2D [30] prioritizes efficiency.
@@
BiDirectional Scan [26] Vision Mamba [26],Motion Mamba [37]
HARMamba [40],MMA [41],VL-Mamba[42]
Video Mamba Suite [43],Point Mamba [44]
LMa-UNet [45]
Motion-Guided Dual-Camera Tracker [46]
Cross-Scan [27] VMamba [27],VL-Mamba[42],VMRNN [47]
RES-VMAMBA [48],Sigma [49],ReMamber [50]
Mamba-UNet [51],Semi-Mamba-UNet [52]
VMambaMorph [53],ChangeMamba [54]
H-vmunet [55],MambaMIR [56],MambaIR [57]
Serpent [58],Mamba-HUNet [59],TM-UNet [60]
Swin-UMamba [61],UltraLight VM-UNet [62]
VM-UNet [63],VM-UNET-V2  [64]
MedMamba [65],MIM-ISTD [66],RS3Mamba [67]
Continuous 2D Scanning [29] PlainMamba [29]
Local Scan [28] LocalMamba [28],FreqMamba [68]
Efficient 2D Scanning (ES2D) [30] EfficientVMamba [30]
Zigzag Scan [34] ZigMa [34]
Omnidirectional Selective Scan [35] VmambaIR [35],RS-Mamba [69]
3D BiDirectional Scan [36] VideoMamba [36]
Hierarchical Scan [37] Motion Mamba [37]
Spatiotemporal Selective Scan [38] Vivim [38]
Multi-Path Scan [39] RSMamba [39]